OpenStudy (anonymous):

Choose the factored form of the given equation. 6x^2 + 2x - 8 = 0

OpenStudy (anonymous):

(6x - 8)(x + 1) = 0 (6x + 1)(x - 8) = 0 (6x + 8)(x - 1) = 0

OpenStudy (anonymous):

This (6x + 8)(x - 1) = 0?

OpenStudy (anonymous):

(6x+8)(x-1)=0

OpenStudy (anonymous):

6x^2-6x+8x-8=0, 6x^2+2x-8=0

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

You're welcome!